OpenStudy (skullpatrol):

The ratio of the sum of two positive integers to their difference is 7:5. If the the sum of the two numbers is at most 25, find all possible values for the pair of numbers.

OpenStudy (rizags):

ok got it

OpenStudy (rizags):

write it out first as such: \[\huge \frac{m+n}{m-n}=\frac{7}{5}\]

OpenStudy (rizags):

Then include this: \[\huge 0<m+n \le25\]

OpenStudy (rizags):

Simplifying the first equation gives \[\huge m=6n\]

OpenStudy (rizags):

find all possible values of that in the interval now

OpenStudy (rizags):

I believe there are 3 sets (a, b)

OpenStudy (anonymous):

We know that \(n\leq m\), because the ratio is positive.

OpenStudy (rizags):

i believe it should be \[\huge n<m\]

OpenStudy (rizags):

but same idea

OpenStudy (rizags):

Hint: try \[\huge n=1, 2, 3....\]

OpenStudy (skullpatrol):

Thanks @Rizags

OpenStudy (rizags):

np